I'm 4o years old. My daughter is 9. How old will I be when I'm 3 times older than she is. I'm working on a formula that looks like this: 3x=x+31. 31 being how old I was at her birth. I'm interested in seeing how this formula is written, as well as how to solve it. Thanks.
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

I'm 4o years old. My daughter is 9. How old will I be when I'm 3 times older than she is. I'm working on a formula that looks like this: 3x=x+31. 31 being how old I was at her birth. I'm interested in seeing how this formula is written, as well as how to solve it.
-------------
The difference will always be 31 (30 or 32 sometimes, depending on bdays)
------
3 times older = 4 times as old.
There's no integer solution for 4 times and 31.
m = mother's age
d = daughter's age
----
m = d + 31
m = 4d
---
4d = d + 31
3d = 31
d = 10 1/3 yrs
m = 41 1/3 yrs
-------
Same for 3 times, no integer solution.
31 is a prime number, so it doesn't work out.
--------
If the difference is, eg, 32 yrs, you can have several multiples with integers.

Answer by jim_thompson5910(35256)

Let x = number of years that pass from now

So in x = 1 year, you will be 40+x = 40+1 = 41 years old and she will be 9+x = 9+1 = 10 years old

So in general, after x years pass, you will be 40+x years old and she will be 9+x years old.


You want to know how long it will take for you to be 3 times as old as her in x years.


So this means that

(Your age in x years) = 3*(Her age in x years)

40+x = 3*(9+x)

So the equation is 40+x = 3*(9+x)

-------------------------------------------------------

Now let's solve 40+x = 3*(9+x)


40+x = 3*(9+x)

40+x = 27+3x

x = 27+3x-40

x-3x = 27-40

-2x = -13

x = -13/(-2)

x = 13/2

x = 6.5

Since this result is not a whole number, this means that you'll never be 3 times as old as her.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------